# Eigenstate solutions of the Fermi-Hubbard model via symmetry-enhanced variational quantum eigensolver

### v1

@@ IV.2.2 Excited states
Figure 4: (a) The two-dimensional Hubbard model with four lattice sites, where circles represent lattice points, the numbers in the middle are the point indices, and arrows indicate that electrons only hop between nearest-neighbor lattice sites. (b) The particle number preserving circuit for the four-site Hubbard model, applying 5 layers for the ground state calculation and 7 layers for the second excited state calculation. (c)-(d) The change of the average fidelity of hardware efficient ansatz and symmetry-preserving ansatz with the number of iterations when solving the ground state. The black vertical dashed line indicates the iteration number when the average fidelity first reaches $\geq 0.95$. All data are the average of 5 independent computations. (e)-(f) Comparison between the two ansatzes when solving the second excited state.
@@
V The partial eigenstate solutions of the four-site 2D Hubbard model
In this section the two-dimensional Hubbard model with four lattice sites, as shown in Fig. 4(a), are considered. Extend the computational processes to a larger system. The arrows represent electron hopping between nearest-neighbor sites, and the term $\sum_{i,j}\sum_{\sigma}\left(\hat{a}_{i\sigma}^{\dagger}\hat{a}_{j\sigma}+\hat{a}_{j\sigma}^{\dagger}\hat{a}_{i\sigma}\right)=\hat{H}_{0}$ in the Hamiltonian is expanded as follows
@@ V The partial eigenstate solutions of the four-site 2D Hubbard model
When to solve the second excited state, the difference becomes even more evident. A 21-layer hardware efficient ansatz without penalty terms (336 parameters and 147 CNOT gates) and a 7-layer symmetry-preserving ansatz with penalty terms (49 parameters and 147 CNOT gates) were employed to solve this eigenstate. The former ansatz achieved an average fidelity of approximately 0.4 at the 500th iteration, whereas the latter ansatz reached a fidelity of about 0.7, as shown in Fig. 4(e)-(f).
@@ VI Conclusion
Based on this foundation, symmetry-enhanced VQE is employed in this paper to solve the energy eigenstates of the one-dimensional Fermi-Hubbard model with two sites and the two-dimensional Fermi-Hubbard model with four sites in this paper, demonstrating the efficiency of symmetry methods in solving the Hubbard model. Three approaches are compared, which are: the hardware efficient ansatz that completely ignores symmetry, the hardware efficient ansatz that partially considers symmetry through the addition of penalty terms, and the symmetry-preserving ansatz. The third approach preserves the particle number of the state during the computation, and a penalty term was added for the $z$-component of the spin operator in the loss function. Our results show that symmetry methods can achieve better outcomes with fewer computational resources when solving the energy eigenstates of the Hubbard model. This suggests that for solving large-scale Hubbard models, symmetries should be analyzed of the system to obtain effective results with reduced analytical effort and numerical computational resources.
@@
All the derivation details, the code used for the simulation program, and the generated data are available at Ref. [39].
@@ Appendix B SO(3) group element represented by SU(2) group parameters
For the SO(2) group parameters $a,b\in\mathbb{C}$, with $\left|a\right|^{2}+\left|b\right|^{2}=1$, the matrix representation of the SO(3) group element in terms of these group parameters is given by
@@
$\beta$ $=$ $-\frac{i}{2}\left(a^{2}-a^{*2}+b^{2}-b^{*2}\right)\,,$
@@
$\delta$ $=$ $\frac{i}{2}\left(a^{2}-a^{*2}-b^{2}+b^{*2}\right)\,,$
@@
$\varepsilon$ $=$ $i\left(a^{*}b^{*}-ab\right)\,,$
@@ Appendix C Derivation steps for operators encoding
$\sigma_{x}\cdot\sigma_{y}$ $=$ $i\sigma_{z}\,,$ (35)
$\sigma_{y}\cdot\sigma_{x}$ $=$ $-i\sigma_{z}\,,$ (36)
@@
$\sigma^{-}$ $=$ $\frac{1}{2}\left(\sigma_{x}+i\sigma_{y}\right)\,,$ (37)
$\sigma^{+}$ $=$ $\frac{1}{2}\left(\sigma_{x}-i\sigma_{y}\right)\,.$ (38)
@@
$+X_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes iY_{j}-iY_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes X_{j}$
$\left.-iY_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes iY_{j}\right)\otimes\prod_{k=j+1}^{2N-1}I_{k}\,.$
@@
$\hat{a}_{i\uparrow}^{\dagger}\hat{a}_{j\uparrow}+\hat{a}_{j\uparrow}^{\dagger}\hat{a}_{i\uparrow}$ $=$ $\frac{1}{2}\prod_{k=0}^{i-1}I_{k}\otimes\left(X_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes X_{j}\right.$
$\left.+Y_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes Y_{j}\right)\otimes\prod_{k=j+1}^{2N-1}I_{k}\,,$
$\hat{a}_{i\downarrow}^{\dagger}\hat{a}_{j\downarrow}+\hat{a}_{j\downarrow}^{\dagger}\hat{a}_{i\downarrow}$ $=$ $\frac{1}{2}\prod_{k=0}^{i-1+N}I_{k}\otimes\left(X_{i+N}\right.$
@@
Appendix D Computational resources, loss functions, initial states, and fidelities for a two-site 1D Hubbard model
@@ D.3 Initial states
The initial states used in the computation process are shown in the table below.

### v2

@@ IV.2.2 Excited states
Figure 4: (a) The one-dimensional Hubbard model with four lattice sites incorporates periodic boundary conditions, where circles represent lattice points, the numbers in the middle are the point indices, and arrows indicate that electrons only hop between nearest-neighbor lattice sites. (b) The particle number preserving circuit for the four-site Hubbard model, applying 5 layers for the ground state calculation and 7 layers for the second excited state calculation. (c)-(d) The change of the average fidelity of hardware efficient ansatz and symmetry-preserving ansatz with the number of iterations when solving the ground state. The black vertical dashed line indicates the iteration number when the average fidelity first reaches $\geq 0.95$. All data are the average of 5 independent computations. (e)-(f) Comparison between the two ansatzes when solving the second excited state.
@@
V The partial eigenstate solutions of the four-site 1D Hubbard model
In this section the one-dimensional Hubbard model with four lattice sites, as shown in Fig. 4(a), are considered. Extend the computational processes to a larger system. This model incorporates periodic boundary conditions, which reduce the impact of boundary effects. The arrows represent electron hopping between nearest-neighbor sites, and the term $\sum_{i,j}\sum_{\sigma}\left(\hat{a}_{i\sigma}^{\dagger}\hat{a}_{j\sigma}+\hat{a}_{j\sigma}^{\dagger}\hat{a}_{i\sigma}\right)=\hat{H}_{0}$ in the Hamiltonian is expanded as follows
@@ V The partial eigenstate solutions of the four-site 1D Hubbard model
When to solve the second excited state, the difference becomes even more evident. A 21-layer hardware efficient ansatz without penalty terms (336 parameters and 147 CNOT gates) and a 7-layer symmetry-preserving ansatz with penalty terms (49 parameters and 147 CNOT gates) were employed to solve this eigenstate. The former ansatz achieved an average fidelity of approximately 0.4 at the 500th iteration, whereas the latter ansatz reached a fidelity of about 0.7, as shown in Fig. 4(e)-(f). The achievable precision is limited by hardware constraints. Although the optimal fidelity was not reached, the comparison mainly serves to demonstrate the advantage of the symmetry ansatz, which is evident from the results.
@@ VI Conclusion
Based on this foundation, symmetry-enhanced VQE is employed in this paper to solve the energy eigenstates of the one-dimensional Fermi-Hubbard model with two and four sites, demonstrating the efficiency of symmetry methods in solving the Hubbard model. Three approaches are compared, which are: the hardware efficient ansatz that completely ignores symmetry, the hardware efficient ansatz that partially considers symmetry through the addition of penalty terms, and the symmetry-preserving ansatz. The third approach preserves the particle number of the state during the computation, and a penalty term was added for the $z$-component of the spin operator in the loss function. Our results show that symmetry methods can achieve better outcomes with fewer computational resources when solving the energy eigenstates of the Hubbard model. This suggests that for solving large-scale Hubbard models, symmetries should be analyzed of the system to obtain effective results with reduced analytical effort and numerical computational resources. In future research, when computing higher-dimensional Hubbard models, we should also consider the symmetry of the system to verify the effectiveness of this method and apply it to more specific physical systems.
@@
All the derivation details, the code used for the simulation program, and the generated data are available at Ref. [ZenodoOpenSource].
@@ Appendix B SO(3) group element represented by SU(2) group parameters
For the SU(2) group parameters $a,b\in\mathbb{C}$, with $\left|a\right|^{2}+\left|b\right|^{2}=1$, the matrix representation of the SO(3) group element in terms of these group parameters is given by
@@
$\beta$ $=$ $-\frac{\mathrm{i}}{2}\left(a^{2}-a^{*2}+b^{2}-b^{*2}\right)\,$,
@@
$\delta$ $=$ $\frac{\mathrm{i}}{2}\left(a^{2}-a^{*2}-b^{2}+b^{*2}\right)\,$,
@@
$\varepsilon$ $=$ $\mathrm{i}\left(a^{*}b^{*}-ab\right)\,$,
@@ Appendix C Derivation steps for operators encoding
$\sigma_{x}\cdot\sigma_{y}$ $=$ $\mathrm{i}\sigma_{z}\,$, (35)
$\sigma_{y}\cdot\sigma_{x}$ $=$ $-\mathrm{i}\sigma_{z}\,$, (36)
@@
$\sigma^{-}$ $=$ $\frac{1}{2}\left(\sigma_{x}+\mathrm{i}\sigma_{y}\right)\,$, (37)
$\sigma^{+}$ $=$ $\frac{1}{2}\left(\sigma_{x}-\mathrm{i}\sigma_{y}\right)\,.$ (38)
@@
$+X_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes\mathrm{i}Y_{j}-\mathrm{i}Y_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes X_{j}$
$\left.-\mathrm{i}Y_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes\mathrm{i}Y_{j}\right)\otimes\prod_{k=j+1}^{2N-1}I_{k}\,.$
@@
$\hat{a}_{i\uparrow}^{\dagger}\hat{a}_{j\uparrow}+\hat{a}_{j\uparrow}^{\dagger}\hat{a}_{i\uparrow}$ $=$ $\frac{1}{2}\prod_{k=0}^{i-1}I_{k}\otimes\left(X_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes X_{j}\right$. (41)
$\left.+Y_{i}\otimes\prod_{k=i+1}^{j-1}Z_{k}\otimes Y_{j}\right)$
$\otimes\prod_{k=j+1}^{2N-1}I_{k}\,$,
$\hat{a}_{i\downarrow}^{\dagger}\hat{a}_{j\downarrow}+\hat{a}_{j\downarrow}^{\dagger}\hat{a}_{i\downarrow}$ $=$ $\frac{1}{2}\prod_{k=0}^{i-1+N}I_{k}\otimes\left(X_{i+N}\right$. (42)
@@
Appendix D Computational resources, loss functions, initial states, and fidelities for the two-site Hubbard model
@@ D.3 Initial states
The initial states used in the computation process are shown in Table 4.